Hall A Experiment #: E00-114
06/14/04 0800-1600
Shift Leader: J. P. Chen
Shift Workers: R. Wilson,, E. Burtin
0800 | 0900 | 1000 | 1100 | 1200 | 1300 | 1400 | 1500 | Total Time | |
---|---|---|---|---|---|---|---|---|---|
Acceptable Beam in Use | 0.00 | 0.00 | 0.00 | 0.33 | 0.67 | 0.75 | 0.75 | 0.75 | 3.25 |
Beam Available but Not in Use | 0.00 | 0.00 | 0.00 | 0.33 | 0.25 | 0.17 | 0.18 | 0.18 | 1.12 |
Beam Not Available or Unacceptable | 1.00 | 1.00 | 1.00 | 0.33 | 0.08 | 0.08 | 0.07 | 0.06 | 3.63 |
Accelerator Con- figuration Change | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Experiment Ready | 1.00 | 1.00 | 1.00 | 0.67 | 0.75 | 0.83 | 0.82 | 0.82 | 6.88 |
Planned Config- uration Change | 0.00 | 0.00 | 0.00 | 0.33 | 0.25 | 0.17 | 0.18 | 0.18 | 1.12 |
Unplanned Experiment Down | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Hall not ready reasons | |||||||||
Target | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Data Acq | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Other | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
HRSL | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
HRSR | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
HALL_OFF | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
Comments: 8:00 no beam, Accerelator recovering from power loss. 9:00 Oppotunistic controlled access for detector, target and spectrometer work. 11:20 back to beam permit -12:25 with new optics checked raster 12:30 With 4He, target density fluctuation tests 13:40 target tempertaure vs cooling power test 14:35 dithering run 15:00 production run while Compton tuning 15:55 production
Data Entry by: J. P. Chen